# Ion counting and temperature determination of Coulomb-crystallized laser-cooled ions in traps using convolutional neural networks

### v1

@@ III Results and Discussion
III.1 Image classification examples
@@
Figure 3: Training loss (in green) and validation accuracy (in blue) for classifying (a) $N$ in the range of 100–299 ions in steps of 1 ion using the CNN model ResNet18, (b) $T$ in the range of 5–15 mK in steps of 1 mK using AlexNet.
@@
For the example of classifying images with $N$=100–299 in steps of 1 ion mentioned above, the training loss and validation accuracy during training are plotted as a function of training epoch in Fig. 3(a), for which the CNN model ResNet18 with its PTW was employed. The training loss refers to the average cross-entropy loss across all batches in the epoch [39, 47], while the validation accuracy refers to the percentage of correctly predicted classes among all predictions made on the validation dataset. During the training loop, the training loss first dropped sharply for the first tens of epochs and then decreased steadily before it stabilized to $\leqslant 0.001$ after about 157 epochs. The validation accuracy shows an opposite trend, increasing from less than 5% to a final stabilized value of 93% after 155 epochs, thus becoming increasingly capable of making correct predictions. It shows clearly that the model’s performance improved significantly during training and could eventually distinguish the validation images with a difference in the number of ions of one ion count with an accuracy better than 90%.
@@
As an example for classifying images of Coulomb crystals with $T=$5–15 mK in steps of  1 mK, we loaded a total of 2200 images for training, i.e., 200 images are assigned to each class of $T$. The pre-trained model AlexNet was utilized in this example. Except that the images were converted to RGB, as explained in Sec. II.3, the same procedures and hyperparameters (such as batch size, learning rate, number of epochs, etc.) as in the case for classifying $N$ were taken for the training and validation process. The training loss and validation accuracy during training are shown in Fig. 3(b). Both the loss and accuracy converged after $\approx$120 epochs, earlier than the case for $N$ in Fig. 3(a). The final accuracy of 92% highlights the model’s capability of differentiating images with a temperature difference of 1 mK.
@@
To evaluate the distribution of errors, defined as the difference between the model’s predicted class label and the true class label, for the trained models, histograms of errors when applying the models to validation images for the two examples are plotted in Fig. 4. Both histograms indicate a narrow error distribution, with variances of (a) 0.0806 ion counts and (b) 0.0815 mK, respectively. The histogram in Fig. 4(a) is slightly more symmetric than that in Fig. 4. This suggests that there are no systematic biases when applying the model to classify $N$ for the example whereas there is a small bias towards lower temperatures in classifying $T$, although the errors are centered closely on zero (with means -0.004 and -0.018, respectively).
@@
Figure 4: (a) Histogram of errors when using the trained model ResNet18 to classify 1440 validation images with respect to $N$ in the range of 100–299 ions. (b) Histogram of errors when using the trained model AlexNet to classify 440 validation images with respect to $T$ in the range of 5–15 mK.
@@
Note that in the above examples, we let the training loop complete a total of 200 epochs for the sake of illustration. However, to avoid possible overfitting after the convergence where the model becomes too specialized in the training data but fails to perform well on new, unseen data [48, 49, 50, 51], we terminated the training process earlier in practice when the validation accuracy ceased to improve for 10 consecutive epochs. The model including its architecture and updated weights was saved at the epoch in which the highest accuracy was achieved.
@@ III.2 CNN model comparison
Acc. 94% 98% 97% 95%
In the above examples, the classification of $N$ and $T$ over certain ranges was demonstrated using the ResNet18 and AlexNet models, respectively, which proved to give the best performances in comparison with the other CNN models we evaluated. Fig. 5 shows the validation accuracies of all the selected models for classifying $N$ in the range of 100–299 ions in steps of 1 ion and $T$ within 5–15 mK in steps of 1 mK. All models for classifying $N$ (or $T$) were trained on the same set of images with the same hyperparameters. In this figure, the validation accuracies of models without loading their respective pre-trained weights (PTW) in the training are also depicted. The classification performances generally decrease for both $N$ and $T$ when no PTW was utilized, which confirms the merit of transfer learning. However, the extent to which the accuracies deteriorate vary by model. For example, the use of PTW makes little difference in performance for the model Vgg16, whereas the considerable differences for the model MnasNet1_0 make the use of PTW necessary in this case.
The figure also shows the importance of model selection in order to optimize the classification performance. We see that not only do different models perform quite differently for a specific task ($N$ or $T$), but the same model can also lead to very different results for different tasks. It also indicates that the model Vgg16 is suitable for both tasks with satisfactory performances.
Note that in our examples, to achieve similarly high validation accuracy (93% and 92% for $N$ and $T$, respectively), almost six times more images (200 vs. 36) are required for each class for the classification of $T$ with 11 classes than for that of $N$ with 200 classes. Indeed, classifying $T$ appeared to be much more challenging than $N$. For example, when we increased the range of $T$ to be classified from 5–15 mK (step: 1) to 5–25 mK (step: 1), the validation accuracy after convergence dropped to around 79% with the AlexNet model under the same training conditions. Although $T$=5-15 mK is typical temperature range of Coulomb crystals in practical situations [36], we found that a wider range of 5–25 mK could be classified to a high accuracy of 98% with ResNet18 if we could allow the step to be 2 mK, meaning that the model can classify images to temperature differences within 2 mK. For a comparison of classifying $T$=5–25 mK (step: 2 mK) with different CNN models, see Table 1.
Figure 5: Validation accuracies after convergence for different CNN models with and without using their corresponding pre-trained weights (PTW) in the training for classifying $N$=100–299 ions (step: 1 ion) and for classifying $T$=5–15 mK (step: 1 mK)
@@ III.3 Application to the analysis of experimental images
Once a CNN model is trained on the simulated images, the weights of the model with the highest performance saved during training can be used in evaluation mode to classify other images of interest, whether they are from simulations or experiments. While applying the model to simulated images can give a high classification accuracy, as suggested in Sec. III.1, its performance on experimental images may be significantly affected by the quality of the images. Experimental images can differ from simulated ones in terms of noise levels, distortions, contrast, etc., and they are susceptible to changes in experimental conditions. These differences pose challenges to the model’s transferability from simulations to experiments.
@@
In spite of these challenges, we find that a few measures can be taken to alleviate the reduction of the model’s performance when implemented to experiments. Firstly, the parameters used for generating the simulated training images should match or cover the experimental ones. The code used for this paper can easily be adapted to generate simulated images under different experimental parameters on which a new model can be trained. Secondly, when the experiment needs not to discern the variations of temperature or number of ions down to a precision of 1, the model can be trained on images with a higher variation. The classification accuracy should in general increase as the targeted precision is reduced, as can be inferred from Sec. III.2. Last but not least, data augmentation [39, 40, 41, 52, 53] as an effective method developed in machine learning and computer vision can be employed before the images are fed into training. It is used to increase the size and diversity of training datasets by applying various transformations to the existing images, which can include rotations, scaling, noise injection, etc. It aims to improve the robustness and generalization ability of the machine-learned models, thus enhancing their performance on new, unseen images.
@@
Taking the experimental image shown in Fig. 6(a) as an example, the AlexNet model trained on simulated images to classify $T$ in the range of 5–25 mK in steps of 2 mK predicted a temperature of 5 mK. Meanwhile, the ResNet18 model trained to classify $N$ in the range of 100–299 ions in larger steps of 10 ions predicted 200 ions. The corresponding simulated image with $T=$ 5 mK and $N=$200 ions is shown in Figure 6(b). Noticing that the experiment contains noise in contrast to the simulations, we performed data augmentation by adding Gaussian noise with a mean $\mu=0$ and a variance $\sigma^{2}=0.1$ grayscale intensity values (which was chosen after testing with a series of variances: 0.01, 0.05, 0.1, 0.15, 0.2, 0.25) to all the simulated images and then retrained the two models. At the cost of a slightly reduced validation accuracy of 86% with AlexNet, the thus obtained model now predicted a temperature of 9 mK for the experimental image. Meanwhile, the retrained ResNet18 model with a validation accuracy of 87% predicted 190 ions for the same image. A direct visual comparison of the corresponding noise-free and noise-augmented simulated images with $T=$ 9 mK and $N=$190 ions shown in Fig. 6(c) and 6(d), respectively, with the experiment in Fig. 6(a) suggests the new predictions to be a more realistic estimate of these attributes compared to the initial guess obtained without adding noise. Two additional examples are given in Fig. 6(e-h), which compares the experimental images with the noise-added simulated images in which $T$ and $N$ are predicted by the retrained models. By visual comparison of the images, we believe the predicted values provide a fair representation of the experiments.
@@
Figure 6: Comparison between experimental and simulated false-color images with $T$s and $N$s predicted by CNN models using the experiment as inputs. (a) First example experimental image. (b) Simulation of (a) with 200 ions at 5 mK. (c) Simulation of (a) with 190 ions at 9 mK. (d) The same simulated image as (c) but with added Gaussian noise. (e) Second example experimental image. (f) Simulation of (e) with 240 ions at 11 mK. (g) Third example experimental image. (h) Simulation of (g) with 160 ions at 13 mK.
@@
Apart from applying the model a posteriori to pre-recorded experimental images, it can also be directly integrated into an experiment’s data acquisition system, allowing an in-situ determination of the relevant Coulomb crystal parameters in real time. According to our evaluation results on a regular lab computer (CPU: Inter Core i7-3930K, RAM: 16 GB, no GPUs), the time for a model to classify an image with 512$\times$512 pixels is a few tens of milliseconds, sufficiently fast to capture changes in crystal characteristics in real time in typical experiments with trapped ions.
@@ IV Conclusion and outlook
We demonstrated a fast, accurate, and general methodology for determining the number and secular temperature of ions in Coulomb crystals utilizing convolutional neural networks (CNNs). This methodology involves the generation of a substantial dataset of images depicting crystals with varying parameters, specifically the number of ions and their temperatures, through molecular dynamics simulations. These images serve as training set for CNN models tasked with classifying the aforementioned parameters. Upon evaluating the classification performance of several widely-used pre-trained CNN models, we identified ResNet18 and AlexNet as the most effective for classifying $N$ and $T$, respectively. Specifically, for crystals with an ion count ranging from 100 to 299 and temperatures between 5 and 15 mK, the best model achieves an accuracy of 93% for the ion count and an accuracy of 92% for the temperature when discerning variations as minute as one ion and 1 mK, respectively.
@@
With a sightly reduced accuracy, the CNN models trained on simulated images can be used for the classification of experimental images. We suggested several ways to improve the performance of experimental implementations, including reducing the classification precision, and pre-processing the simulated input images (e.g. adding noise) before the training. With a direct integration of the model into the experiments, real-time and in-situ determination of crystal parameters can now be feasible, thereby significantly enhancing the monitoring and control of trapped ions in such experiments.
@@
We note that due to the limited number of experimental images with clearly defined ion numbers and temperatures available for model evaluation, the actual performance of the models requires careful verification in real-life experimental settings. Clearly, the quality of the experimental image to be analyzed plays a critical role. Additionally, the manual classification of experimental images by visual comparison with simulated ones which are used to validate the predictions from the models may be subject to inherent inaccuracies. Ideally, the models’ predictions would be more reliable if they were trained on a large dataset of experimental images labeled using independent ion counting and thermometry methods, as mentioned in Sec. I. When an insufficient number of experimental images is available, incorporating simulated images into a ”hybrid” experimental/ simulated training data set could be beneficial. However, if the models are trained partially or entirely on simulated images, potential systematic errors should be carefully calibrated before applying them to experiments.
@@
The present method is computationally efficient and enables fast data preparation, model training, and evaluation, while also being easily adaptable to different experimental parameters as needed. Furthermore, it can be readily adapted for other ion-trapping experiments, thereby enhancing the ion-counting and thermometry toolkit for trapped ions in the intermediate temperature range. This compares with other ongoing work of training a general model designed to accommodate a broader range of experimental conditions [54]. Potentially it can also be extended to Coulomb crystals with mixed ions, facilitating studies with sympathetically cooled ions [7, 55] (and the references therein) and dynamics of ion-neutral collisions [7, 56, 57, 58, 59, 60].
Code availability
The code for generating simulated images of Coulomb crystals, training and evaluation of CNN models for ion counting and temperature determination of crystals can be found on the GitHub repository with URL: https://github.com/yanningyin/Coulomb-crystal-CNN or on the Zenodo repository with the identifier doi: 10.5281/zenodo.14926184.

### v2

@@ III Results and Discussion
III.1 Image-classification examples
@@
Figure 3: Training loss (in green) and validation accuracy (in blue) for classifying (a) NN in the range of 100–299 ions in steps of 1 ion using the CNN model ResNet18, (b) $T$ in the range of 5–15 mK in steps of 1 mK using AlexNet.
To demonstrate the performance of the CNN models, we present examples of image classification with $N$ in the range 100–299 and $T$ between 5 and 15 mK. This range is representative of typical ion-trap experiments involving medium-sized Coulomb crystals under Doppler cooling. Focusing on these conditions allows us to benchmark the method against realistic experimental settings and establishes a foundation for extention to broader regimes. In principle, the approach is scalable to other $N$ and $T$ ranges, as it primarily requires training models on newly simulated (or experimentally collected) images under the relevant conditions.
@@
For the example of classifying images with NN=100–299 in steps of 1 ion mentioned above, the training loss and validation accuracy during training are plotted as a function of training epoch in Fig. 3(a), for which the CNN model ResNet18 with its PTWs was employed. The training loss refers to the average cross-entropy loss across all batches in the epoch [39, 49], while the validation accuracy refers to the percentage of correctly predicted classes among all predictions made on the validation dataset. During the training loop, the training loss first dropped sharply for the first tens of epochs and then decreased steadily before it stabilized to $\leqslant 0.001$ after about 157 epochs. The validation accuracy showed an opposite trend, increasing from less than 5% to a final stabilized value of 93% after 155 epochs, thus becoming increasingly capable of making correct predictions. This clearly shows that the model’s performance improved significantly during training and could eventually distinguish validation images differing by a single ion with an accuracy better than 90%.
@@
As an example for classifying images of Coulomb crystals with T=T=5–15 mK in steps of  1 mK, we loaded a total of 2200 images for training, i.e., 200 images are assigned to each $T$ class. The pretrained model AlexNet was used in this example. Apart from the fact that the images were converted to RGB, as explained in Sec. II.3, the same procedures and hyperparameters (such as batch size, learning rate, number of epochs, etc.) as in the case for classifying $N$ were taken for the training and validation processes. The training loss and validation accuracy during training are shown in Fig. 3(b). Both the loss and accuracy converged after $\approx$120 epochs, earlier than the case for $N$ in Fig. 3(a). The final accuracy of 92% highlights the model’s ability to differentiate images with a temperature difference of 1 mK.
@@
To evaluate the distribution of errors, defined as the difference between the model’s predicted class label and the true class label, for the trained models, histograms of errors when applying the models to validation images for the two examples are plotted in Figs. 4(a) and 4(b). Both histograms indicate a narrow error distribution, with variances of 0.0806 ion counts and 0.0815 mK [Figs. 4(a) and 4(b), respectively]. The histogram in Fig. 4(a) is slightly more symmetric than that in Fig. 4(b). This suggests that there are no systematic biases when applying the model to classify $N$ for the example, whereas there is a small bias toward lower temperatures in classifying $T$, although the errors are centered closely on zero (with means of -0.004 and -0.018, respectively). Figures 4(c) and 4(d) show the models’ predictions for $N$ and $T$ versus the true labels of the input images, indicating both high accuracy and high precision.
@@
Figure 4: (a) Histogram of errors when using the trained model ResNet18 to classify 1440 validation images with respect to NN in the range 100–299 ions. (b) Histogram of errors when using the trained model AlexNet to classify 440 validation images with respect to $T$ in the range 5–15 mK. (c) Predicted NN by the model vs the true NN of the input images. Only a proportion of the datapoints in the range NN=100–299 ions are shown for clarity. (d) Predicted $T$ vs true $T$.
@@
Note that in the above examples, we let the training loop complete a total of 200 epochs for the sake of illustration. However, to avoid possible overfitting after the convergence—whereby the model becomes too specialized in the training data but fails to perform well on new, unseen data [50, 51, 52, 53]—in practice, we terminated the training process earlier at the point when the validation accuracy ceased to improve for ten consecutive epochs. The model, including its architecture and updated weights, was saved at the epoch in which the highest accuracy was achieved.
@@ III.2 CNN model comparison
Accuracy 94% 98% 97% 95%
In the above examples, the classification of $N$ and $T$ over certain ranges was demonstrated using the ResNet18 and AlexNet models, respectively, which were found to give the best performances in comparison with the other CNN models we evaluated. Figure 5 shows the validation accuracies of all the selected models for classifying $N$ in the range 100–299 ions in steps of 1 ion and $T$ within 5–15 mK in steps of 1 mK. All models for classifying $N$ (or $T$) were trained on the same set of images with the same hyperparameters. In this figure, the validation accuracies of models without loading their respective PTWs in the training are also depicted. The classification performances generally decreased for both $N$ and $T$ when no PTWs were used, which confirms the merit of transfer learning. However, the extent to which the accuracies deteriorate vary by model. For example, the use of PTWs makes little difference in the performance of the VGG16 model, suggesting that it can still learn useful patterns from the training data even when starting from random weights. In contrast, the MnasNet1_0 model showed a large drop in accuracy without PTWs. This is likely because MnasNet1_0 is a smaller and more compact model designed to run efficiently on mobile or low-power devices. As a result, it has less capacity to learn complex patterns from scratch and therefore benefits more from being initialized with knowledge learned from a large dataset. For such models, starting with PTWs is especially important to achieve good performance.
Figure 5 also shows the importance of model selection to optimize the classification performance. We see that not only do different models perform quite differently for a specific task ($N$ or $T$), but the same model can also lead to very different results for different tasks. This also suggests that the VGG16 model is suitable for both tasks, achieving satisfactory performance.
Note that in our examples, to achieve similarly high validation accuracy (93% and 92% for $N$ and $T$, respectively), almost six times more images (200 vs. 36) are required for each class for the classification of $T$ with 11 classes than for that of $N$ with 200 classes. Indeed, classifying $T$ appeared to be much more challenging than classifying $N$. For example, when we increased the range of $T$ to be classified from 5–15 mK (step: 1 mK) to 5–25 mK (step: 1 mK), the validation accuracy after convergence dropped to around 79% with the AlexNet model under the same training conditions. Although TT=5-15 mK is a typical temperature range of Coulomb crystals in practical situations [36], we found that a wider range of 5–25 mK could be classified to a high accuracy of 98% with ResNet18 if we could allow the step to be 2 mK, meaning that the model can classify images to temperature differences within 2 mK. For a comparison of classifying TT=5–25 mK (step: 2 mK) with different CNN models, see Table 1.
Figure 5: Validation accuracies after convergence for different CNN models with and without using their corresponding pretrained weights (PTWs) in the training for classifying NN=100–299 ions (step: 1 ion) and for classifying TT=5–15 mK (step: 1 mK)
@@ III.3.1 Transferring from simulations to experiments
Once a CNN model is trained on simulated images, the weights of the model with the highest performance saved during training can be used in evaluation mode to classify other images of interest, whether they are from simulations or experiments. Although applying the model to simulated images can give a high classification accuracy, as suggested in Sec. III.1, its performance on experimental images may be significantly affected by the quality of the images. Experimental images can differ from simulated ones in terms of noise levels, distortions, contrast, etc., and they are susceptible to changes in experimental conditions. These differences pose challenges to the model’s transferability from simulations to experiments.
@@
In spite of these challenges, we found that a few measures can be taken to alleviate the reduction of the model’s performance when applied to experiments. First, the parameters used for generating the simulated training images should match or cover the experimental ones. The code used for this paper can easily be adapted to generate simulated images under different experimental parameters on which a new model can be trained. Second, when the experiment does not need to discern the variations of temperature or number of ions down to a precision of 1, the model can be trained on images with a higher variation. The classification accuracy should in general increase as the targeted precision is reduced, as can be inferred from Sec. III.2. Last but not least, data augmentation [39, 40, 41, 54, 55] as an effective method developed in machine learning and computer vision can be employed before the images are fed into training. It can be used to increase the size and diversity of training datasets by applying various transformations to existing images, which can include rotations, scaling, noise injection, etc. This approach aims to improve the robustness and generalization ability of the machine-learned models, thus enhancing their performance on new, unseen images.
@@
Taking the experimental image shown in Fig. 6(a) as an example, the AlexNet model trained on simulated images to classify $T$ in the range 5–25 mK in steps of 2 mK predicted a temperature of 5 mK. Meanwhile, the ResNet18 model trained to classify $N$ in the range 100–299 ions in larger steps of 10 ions predicted 200 ions. The corresponding simulated image with $T=5$ mK and $N=200$ ions is shown in Fig. 6(b). Noticing that, in contrast to the simulated images, the experimental images contains noise, we performed data augmentation by adding Gaussian noise with mean $\mu=0$ and variance $\sigma^{2}=0.1$ grayscale intensity values (which was chosen after testing with a series of variances: 0.01, 0.05, 0.1, 0.15, 0.2, 0.25) to all the simulated images and then retrained the two models. This common data-augmentation strategy helps CNNs generalize better by narrowing the gap between clean simulated inputs and noisy experimental data. At the cost of a slightly reduced validation accuracy of 86% with AlexNet, the thus-obtained model now predicted a temperature of 9 mK for the experimental image. Meanwhile, the retrained ResNet18 model with a validation accuracy of 87% predicted 190 ions for the same image. A direct visual comparison of the corresponding noise-free and noise-augmented simulated images with $T=$ 9 mK and N=N=190 ions shown in Fig. 6(c) and 6(d), respectively, with the experiment in Fig. 6(a), suggests that the new predictions provide a more realistic estimate of these attributes compared to the initial guess obtained without adding noise. Two additional examples are given in Fig. 6(e-h), which compare the experimental images with the noise-added simulated images in which $T$ and $N$ are predicted by the retrained models. By visual comparison, we believe that the predicted values provide a fair representation of the experiments.
@@
Figure 6: Comparison between experimental and simulated false-color images with $T$ and $N$ values predicted by CNN models using the experiment as inputs. (a) First example experimental image. (b) Simulation of (a) with 200 ions at 5 mK. (c) Simulation of (a) with 190 ions at 9 mK. (d) The same simulated image as (c) but with added Gaussian noise. (e) Second example experimental image. (f) Simulation of (e) with 240 ions at 11 mK. (g) Third example experimental image. (h) Simulation of (g) with 160 ions at 13 mK.
III.3.2 Models trained directly on experimental images
While we were unable to directly verify the output of the CNN models for large crystals with independent ion counting and thermometry methods (other than by visual comparison), we evaluated the generalizability and reliability of the present CNN approach using experimental images of Coulomb crystals containing small numbers of ions. Experimentally, we prepared a one-dimensional (1D) crystal with a defined number of ions (up to 12), and then transformed it to a 3D crystal by adjusting our trap voltages. To ensure that no ions were lost during the transformation, we repeatedly reverted the system to the 1D-crystal configuration before restoring it to the 3D case. Images of both cases with known labels of $N$ were saved and then used to train a CNN model to infer $N$.
We collected a total of 70 experimental images across seven classes of $N$ (4, 6, 7, 9, 10, 11, and 12), with each class containing five images of 1D crystals and five images of 3D crystals. Example images are shown in Fig. 7(a). After training a ResNet18 model with the same configurations as described in Sec. II on these images, the validation accuracy stabilized at 100% after around 40 epochs. Figure 7(b) shows the evolution of the training loss and validation accuracy.
Given the small dataset size (70 images across seven classes), where random guessing would yield an accuracy of only $1/7\approx 14.3$%, the model’s perfect accuracy demonstrates a strong learning capability even with limited real-world data, highlighting the extensibility and reliability of the approach.
Figure 7: (a) Example experimental images (false color) of 1D (left) and 3D (right) Coulomb crystals with different NN. (b) Training loss (in green) and validation accuracy (in blue) during training directly on experimental images.
III.3.3 Potential for in-situ and real-time inference
@@
In addition to applying the model a posteriori to prerecorded experimental images, it can be directly integrated into an experiment’s data-acquisition system, allowing an in-situ determination of the relevant Coulomb-crystal parameters in real time. According to our evaluation results on a regular laboratory computer (CPU: Intel Core i7-3930K, RAM: 16 GB, no GPUs), the time taken for a model to classify an image with $\times512$ pixels is a few tens of milliseconds, sufficiently fast to capture changes in crystal characteristics in real time in typical experiments with trapped ions.
@@ IV Conclusion and outlook
We demonstrated a fast, accurate, and general methodology for determining the number and secular temperature of ions in Coulomb crystals using CNNs. This methodology involves the generation of a substantial dataset of images depicting crystals with varying parameters, specifically, the number of ions and their temperatures, through molecular-dynamics simulations. These images serve as a training set for CNN models tasked with classifying the aforementioned parameters. Upon evaluating the classification performance of several widely used pretrained CNN models, we identified ResNet18 and AlexNet as the most effective for classifying $N$ and $T$, respectively. Specifically, for crystals with an ion count ranging from 100 to 299 and temperatures between 5 and 15 mK, the best model achieves an accuracy of 93% for the ion count and an accuracy of 92% for the temperature when discerning variations as minute as one ion and 1 mK, respectively.
@@
With a slightly reduced accuracy, the CNN models trained on simulated images can be used for the classification of experimental images. We suggested several ways to improve the performance of experimental implementations, including reducing the classification precision and prepossessing the simulated input images (e.g., adding noise) before the training. Additionally, we demonstrated that CNN models trained directly on a small set of labeled experimental images could achieve high classification accuracy, highlighting the reliability and potential of the present approach. With a direct integration of the model into the experiments, real-time and in-situ determination of crystal parameters is now feasible, thereby significantly enhancing the monitoring and control of trapped ions in such experiments.
@@
We note that due to the limited number of experimental images with clearly defined numbers of ions and temperatures available for model evaluation, the actual performance of the models requires careful verification in real-life experimental settings. Clearly, the quality of the experimental image to be analyzed plays a critical role. Additionally, the manual classification of experimental images by visual comparison with simulated images, which are used to validate the model predictions, may be subject to inherent inaccuracies. Ideally, the models’ predictions would be more reliable if they were trained on a large dataset of experimental images labeled using independent ion counting and thermometry methods, as mentioned in Sec. I. When an insufficient number of experimental images is available, incorporating simulated images into a “hybrid” experimental and simulated training dataset could be beneficial. However, if the models are trained partially or entirely on simulated images, potential systematic errors should be carefully calibrated before applying them to experiments.
@@
The present method is computationally efficient and enables fast data preparation, model training, and evaluation, while also being easily adaptable to different experimental parameters as needed. Furthermore, it can be readily adapted to other ion-trapping experiments, thereby enhancing the ion-counting and thermometry toolkit for trapped ions in the intermediate-temperature range. This compares with other parallel work on training a general model designed to accommodate a broader range of experimental conditions [56]. Potentially, it can also be extended to Coulomb crystals with mixed ions, facilitating studies with sympathetically cooled ions (see Refs. [7, 57] and references therein) and the dynamics of ion-neutral collisions [7, 58, 59, 60, 61, 62].